# Assistance-Seeking in Human-Supervised Autonomy:

### v1

@@ IV-A Expected Reward and Certainty-equivalent Dynamics
Given the probability of successful autonomous collection by the robot, $p^{\textup{suc}}_{C^{1}}$, the probabilities of human action $\mathbb{P}(a^{H}|T,P,C^{1}),\ a^{H}\in\{a^{H+},a^{H-}\}$, for $C^{1}\in\{C^{\textup{1,L}},C^{\textup{1,H}}\}$, and the probability of an object-collection trial being high complexity, $\beta_{1}$, the expected rewards for action $a^{R+}$ in the supervisory task is
${\mathcal{J}}^{\textup{coll}}({T_{t},P_{t},a^{R+}_{t}})$ $=(1-\beta_{1}){\mathcal{R}}^{\textup{coll}}_{C^{1,L}}(T_{t},P_{t},a^{R+}_{t})$
$+\beta_{1}{\mathcal{R}}^{\textup{coll}}_{C^{1,H}}(T_{t},P_{t},a^{R+}_{t}),$ (6)
@@
${\mathcal{R}}^{\textup{coll}}_{C^{1}}(T_{t},P_{t},a^{R+}_{t})$ $=\mathbb{P}(a^{H+}|T_{t},P_{t},C^{1}_{t}){\mathcal{R}}^{\textup{auto-coll}}_{C^{1}}(a^{R+},a^{H+})$
$+\mathbb{P}(a^{H-}|T_{t},P_{t},C^{1}_{t}){\mathcal{R}}^{\textup{coll}}(a^{R+},a^{H-},*),$
@@
Similarly, for the probability of the target-tracking task being normal speed, $\beta_{2}$, the expected reward ${\mathcal{J}}^{\textup{track}}({P_{t},a^{R}_{t}})$ in the target-tracking task is
${\mathcal{J}}^{\textup{track}}({P_{t},a^{R}_{t}})$ $=(1-\beta_{2}){\mathcal{R}}^{\textup{track}}(C^{\textup{2,slow}},p_{t+1})$
@@
$\mathbb{P}(\theta^{1}_{t})$ $=(1-\beta_{1})p^{\textup{suc}}_{C^{\textup{1,L}}}\mathbb{P}(a^{H+}|T,P,C^{1}),$
$\mathbb{P}(\theta^{2}_{t})$ $=(1-\beta_{1})(1-p^{\textup{suc}}_{C^{\textup{1,L}}})\mathbb{P}(a^{H+}|T,P,C^{1}),$
@@
$\mathbb{P}(\theta^{4}_{t})$ $=\beta_{1}p^{\textup{suc}}_{C^{\textup{1,L}}}\mathbb{P}(a^{H+}|T,P,C^{1}),$
$\mathbb{P}(\theta^{5}_{t})$ $=\beta_{1}(1-p^{\textup{suc}}_{C^{\textup{1,L}}})\mathbb{P}(a^{H}|T,P,C^{1}),$
@@
$\mathbb{P}(\theta^{7}_{t})$ $=\mathbb{P}(a^{H-}|T,P,C^{1}).$
@@
$\mathbb{P}(\phi^{1}_{t})$ $=(1-\beta_{2})\epsilon_{t},$
$\mathbb{P}(\phi^{2}_{t})$ $=\beta_{2}\epsilon_{t},$
$\mathbb{P}(\phi^{3}_{t})$ $=(1-\beta_{2})(1-\epsilon_{t}),$
$\mathbb{P}(\phi^{4}_{t})$ $=\beta_{2}(1-\epsilon_{t}),$
$\mathbb{P}(\phi^{5}_{t})$ $=(1-\beta_{2})\epsilon_{t},$
$\mathbb{P}(\phi^{6}_{t})$ $=\beta_{2}\epsilon_{t},$
$\mathbb{P}(\phi^{7}_{t})$ $=(1-\beta_{2})(1-\epsilon_{t}),$
$\mathbb{P}(\phi^{8}_{t})$ $=\beta_{2}(1-\epsilon_{t}),$
@@
$\epsilon_{t}=q_{t-1}\mathbb{E}_{C^{1}_{t-1}\sim\text{Ber}(\beta_{1})}\big{[}p^{\textup{suc}}_{C^{1}}\mathbb{P}(a^{H+}|T_{t-1},P_{t-1},C^{1}_{t-1}))\big{]}\\ +(1-q_{t-1}),$
@@
Given the probability of successful autonomous collection by the robot, $p^{\textup{suc}}_{C^{1}}$, the probabilities of human action $\mathbb{P}(a^{H}|T,P,C^{1}),\ a^{H}\in\{a^{H+},a^{H-}\}$, for $C^{1}\in\{C^{\textup{1,L}},C^{\textup{1,H}}\}$, the certainty-equivalent evolution of the trust $\bar{T}_{t}$ and engagement $\bar{P}_{t}$ dynamics and associated probability of intervention ¯$\bar{\mathbb{P}}_{t}$ and tracking performance $\bar{p}_{t}$ are
@@
$\bar{P}_{t+1}$ $={A}^{P}\bar{P}_{t}+\hat{B}^{P}_{t}u_{t}$
¯$\bar{\mathbb{P}}_{t}$ $=\mathbb{E}_{C^{1}\sim\text{Ber}(\beta_{1})}\big{[}\mathbb{P}(a^{H+}|\bar{T}_{t},\bar{P}_{t},C^{1})\big{]}$
$\bar{p}_{t}$ $=C^{P}\bar{P}_{t},$
@@
$\hat{B}^{P}_{t}$ $=\begin{bmatrix}\sum_{i=5}^{8}\mathbb{P}(\phi^{i}_{t})(B^{P})_{1,i}&\sum_{i=1}^{4}\mathbb{P}(\phi^{i}_{t})(B^{P})_{1,i}\end{bmatrix}$
@@
The entries in $\hat{B}^{T}_{t}$ and $\hat{B}^{P}_{t}$ correspond to the expected influence of the autonomous collection attempt and assistance-seeking on trust and engagement, respectively. They are obtained by averaging entries of $B^{T}$ and $B^{P}$, respectively, associated with these events weighted by their probabilities. ∎
@@ IV-B MPC-based Assistance-seeking Policy
Using the models described in Section III, we design an optimal assistance-seeking policy using Model Predictive Control (MPC) [31]. We adopt a certainty-equivalent MPC approach [32] that considers the expected reward and deterministic system dynamics described in Lemmas 1 and 2, respectively. Specifically, the following optimization problem is solved at the beginning of each trial $t$ with trust and target-tracking engagement $T_{t}$ and $P_{t}$, respectively.
$\begin{split}\underset{{q_{t}},\ldots{q_{t+N-1}}}{\text{maximize}}\quad&\sum_{\tau=t}^{t+N-1}\bar{\mathcal{R}}^{\textup{coll}}({\bar{T}_{\tau},\bar{P}_{\tau},q_{\tau}})+\bar{\mathcal{R}}^{\textup{track}}({\bar{P}_{\tau},q_{\tau}})\\ \text{s.t.}\quad&\bar{T}_{\tau+1}={A}^{T}\bar{T}_{\tau}+\hat{B}^{T}_{\tau}u_{\tau},\ \bar{T}_{t}=T_{t}\\ &\bar{P}_{\tau+1}={A}^{P}\bar{P}_{\tau}+\hat{B}^{P}_{\tau}u_{\tau}\ ,\bar{P}_{t}=P_{t}\\ &\bar{\mathbb{P}}_{T}=\mathbb{E}_{C^{1}\sim\text{Ber}(\beta_{1})}\big{[}\mathbb{P}(a^{H+}|\bar{T}_{t},\bar{P}_{t},C^{1})\big{]}\\ &\bar{p}_{t}=C^{P}\bar{P}_{t}\\ &q_{\tau}\in[0,1],\end{split}$ (7)
@@
¯$\bar{\mathcal{R}}^{\textup{coll}}({T_{\tau},P_{\tau},a^{R}_{\tau}})$ $=q_{\tau}{\mathcal{J}}^{\textup{coll}}_{C^{1}}(T_{\tau},P_{\tau},a^{R+})$
@@
¯$\bar{\mathcal{R}}^{\textup{track}}({P_{\tau},q_{\tau}})$ $=q_{\tau}{\mathcal{J}}^{\textup{track}}({P_{\tau},a^{R+}_{\tau}})$
$+(1-q_{\tau}){\mathcal{J}}^{\textup{track}}({P_{\tau},a^{R-}_{\tau}}).$
@@
In high-complexity object collection tasks, the optimal action is to ask for assistance if secondary task engagement $P_{t}$ falls below a threshold of approximately $P_{t}=5$, regardless of tracking task complexity, previous experience, or trust level. Since low target-tracking performance may be caused by the human paying more attention to the supervisory task, the optimal policy seeks assistance at low target-tracking engagement; thereby allowing the human to finish the tracking task and then assist the manipulator. In contrast, when tracking task engagement is high, the optimal policy depends on human trust: it asks for assistance at low trust levels and collects autonomously for high trust levels. The threshold of trust for seeking assistance decreases with target-tracking performance. Since higher target-tracking performance is an outcome of higher engagement with the tracking task resulting in fewer interruptions by the supervisor, the optimal policy attempts autonomous collection of the object.
@@ IV-C Evaluation Experiment
Each participant completed two blocks of experiments, during which the robot followed the MPC and greedy policies. The order of the two blocks was randomized. In each block, participants engage in 15 trials each of low-complexity and high-complexity supervisory object collection tasks. Half of these trials involved simultaneous execution of slow-speed target tracking tasks, while the remaining involved normal-speed target tracking. Sequences of object collection and tracking tasks with the above-mentioned complexities were selected and these sequences were randomly permuted for every participant.
@@
To match a realistic scenario in which the participants may not report their trust after each trial, we did not use the reported trust values and estimated trust and tracking task engagement of the human participant using their actions to rely or intervene, and the measured tracking performance $p_{t}$.
To this end, we adopted a particle filter with human trust dynamics (1) and engagement dynamics (3), and using the human action model (5) and target-tracking task performance output (4) to estimate $T_{t+1}$ and $P_{t}$.
@@
The cumulative reward for all participants under both the greedy and the MPC policy is illustrated in Fig. 5. The score of a trial is calculated as the sum of rewards in both tasks. The median scores are 65.75 and 57 for the MPC and the greedy policy, respectively. The MPC policy outperforms the greedy policy for most participants. The number of interruptions in our MPC policy, 16, is also lesser than the greedy policy, 23.
@@
Figure 5: Cumulative reward statistics for both policies
V Conclusions

### v2

@@ IV-A Expected Reward and Certainty-equivalent Dynamics
Given the probability of successful autonomous collection by the robot, $p^{\textup{suc}}_{C^{1}}$, the probabilities of human action $\mathbb{P}(a^{H}|T,G,C^{1}),\ a^{H}\in\{a^{H+},a^{H-}\}$, for $C^{1}\in\{C^{\textup{1,L}},C^{\textup{1,H}}\}$, and the probability of an object-collection trial being high complexity, $\beta_{1}$, the expected rewards for action $a^{R+}$ in the supervisory task is
${\mathcal{J}}^{\textup{coll}}({T_{t},G_{t},a^{R+}_{t}})$ $=(1-\beta_{1}){\mathcal{R}}^{\textup{coll}}_{C^{1,L}}(T_{t},G_{t},a^{R+}_{t})$
$+\beta_{1}{\mathcal{R}}^{\textup{coll}}_{C^{1,H}}(T_{t},G_{t},a^{R+}_{t}),$ (6)
@@
${\mathcal{R}}^{\textup{coll}}_{C^{1}}(T_{t},G_{t},a^{R+}_{t})$ $=\mathbb{P}(a^{H+}|T_{t},G_{t},C^{1}_{t}){\mathcal{R}}^{\textup{auto-coll}}_{C^{1}}(a^{R+},a^{H+})$
$+\mathbb{P}(a^{H-}|T_{t},G_{t},C^{1}_{t}){\mathcal{R}}^{\textup{coll}}(a^{R+},a^{H-},*),$
@@
Similarly, for the probability of the target-tracking task being normal speed, $\beta_{2}$, the expected reward ${\mathcal{J}}^{\textup{track}}({G_{t},a^{R}_{t}})$ in the target-tracking task is
${\mathcal{J}}^{\textup{track}}({G_{t},a^{R}_{t}})$ $=(1-\beta_{2}){\mathcal{R}}^{\textup{track}}(C^{\textup{2,slow}},p_{t+1})$
@@
$\mathbb{P}(\theta^{1}_{t})$ $=(1-\beta_{1})p^{\textup{suc}}_{C^{\textup{1,L}}}\mathbb{P}(a^{H+}|T,G,C^{1}),$
$\mathbb{P}(\theta^{2}_{t})$ $=(1-\beta_{1})(1-p^{\textup{suc}}_{C^{\textup{1,L}}})\mathbb{P}(a^{H+}|T,G,C^{1}),$
@@
$\mathbb{P}(\theta^{4}_{t})$ $=\beta_{1}p^{\textup{suc}}_{C^{\textup{1,L}}}\mathbb{P}(a^{H+}|T,G,C^{1}),$
$\mathbb{P}(\theta^{5}_{t})$ $=\beta_{1}(1-p^{\textup{suc}}_{C^{\textup{1,L}}})\mathbb{P}(a^{H}|T,G,C^{1}),$
@@
$\mathbb{P}(\theta^{7}_{t})$ $=\mathbb{P}(a^{H-}|T,G,C^{1}).$
@@
$\mathbb{P}(\phi^{1}_{t})=(1-\beta_{2})\epsilon_{t},$ $\mathbb{P}(\phi^{5}_{t})=(1-\beta_{2})\epsilon_{t},$
$\mathbb{P}(\phi^{2}_{t})=\beta_{2}\epsilon_{t},$ $\mathbb{P}(\phi^{6}_{t})=\beta_{2}\epsilon_{t},$
$\mathbb{P}(\phi^{3}_{t})=(1-\beta_{2})(1-\epsilon_{t}),$ $\mathbb{P}(\phi^{7}_{t})=(1-\beta_{2})(1-\epsilon_{t}),$
$\mathbb{P}(\phi^{4}_{t})=\beta_{2}(1-\epsilon_{t}),$ $\mathbb{P}(\phi^{8}_{t})=\beta_{2}(1-\epsilon_{t}),$
@@
$\epsilon_{t}=q_{t-1}\mathbb{E}_{C^{1}_{t-1}\sim\text{Ber}(\beta_{1})}\big{[}p^{\textup{suc}}_{C^{1}}\mathbb{P}(a^{H+}|T_{t-1},G_{t-1},C^{1}_{t-1}))\big{]}\\ +(1-q_{t-1}),$
@@
Given the probability of successful autonomous collection by the robot, $p^{\textup{suc}}_{C^{1}}$, the probabilities of human action $\mathbb{P}(a^{H}|T,G,C^{1}),\ a^{H}\in\{a^{H+},a^{H-}\}$, for $C^{1}\in\{C^{\textup{1,L}},C^{\textup{1,H}}\}$, the certainty-equivalent evolution of the trust $\bar{T}_{t}$ and engagement $\bar{G}_{t}$ dynamics and associated probability of intervention ¯$\bar{\mathbb{P}}_{t}$ and tracking performance $\bar{p}_{t}$ are
@@
$\bar{G}_{t+1}$ $={A}^{G}\bar{G}_{t}+\hat{B}^{G}_{t}u_{t}$
¯$\bar{\mathbb{P}}_{t}$ $=\mathbb{E}_{C^{1}\sim\text{Ber}(\beta_{1})}\big{[}\mathbb{P}(a^{H+}|\bar{T}_{t},\bar{G}_{t},C^{1})\big{]}$
$\bar{p}_{t}$ $=C^{G}\bar{G}_{t},$
@@
$\hat{B}^{G}_{t}$ $=\begin{bmatrix}\sum_{i=5}^{8}\mathbb{P}(\phi^{i}_{t})(B^{G})_{1,i}&\sum_{i=1}^{4}\mathbb{P}(\phi^{i}_{t})(B^{G})_{1,i}\end{bmatrix}.$
@@
The entries in $\hat{B}^{T}_{t}$ and $\hat{B}^{G}_{t}$ correspond to the expected influence of the autonomous collection attempt and assistance-seeking on trust and engagement, respectively. They are obtained by averaging entries of $B^{T}$ and $B^{G}$, respectively, associated with these events weighted by their probabilities. ∎
@@ IV-B MPC-based Assistance-seeking Policy
Using the models described in Section III, we design an optimal assistance-seeking policy using Model Predictive Control (MPC) [31]. We adopt a certainty-equivalent MPC approach [32] that considers the expected reward and deterministic system dynamics described in Lemmas 1 and 2, respectively. Specifically, the following optimization problem is solved at the beginning of each trial $t$ with trust and target-tracking engagement $T_{t}$ and $G_{t}$, respectively.
$\begin{split}\underset{{q_{t}},\ldots{q_{t+N-1}}}{\text{maximize}}\quad&\sum_{\tau=t}^{t+N-1}\bar{\mathcal{R}}^{\textup{coll}}({\bar{T}_{\tau},\bar{G}_{\tau},q_{\tau}})+\bar{\mathcal{R}}^{\textup{track}}({\bar{G}\tau,q_{\tau}})\\ \text{s.t.}\quad&\bar{T}_{\tau+1}={A}^{T}\bar{T}_{\tau}+\hat{B}^{T}_{\tau}u_{\tau},\ \bar{T}_{t}=T_{t}\\ &\bar{G}_{\tau+1}={A}^{G}\bar{G}_{\tau}+\hat{B}^{G}_{\tau}u_{\tau}\ ,\bar{G}_{t}=G_{t}\\ &\bar{\mathbb{P}}_{T}=\mathbb{E}_{C^{1}\sim\text{Ber}(\beta_{1})}\big{[}\mathbb{P}(a^{H+}|\bar{T}_{t},\bar{G}_{t},C^{1})\big{]}\\ &\bar{p}_{t}=C^{G}\bar{G}_{t}\\ &q_{\tau}\in[0,1],\end{split}$ (7)
@@
¯$\bar{\mathcal{R}}^{\textup{coll}}({T_{\tau},G_{\tau},a^{R}_{\tau}})$ $=q_{\tau}{\mathcal{J}}^{\textup{coll}}_{C^{1}}(T_{\tau},G_{\tau},a^{R+})$
@@
¯$\bar{\mathcal{R}}^{\textup{track}}({G_{\tau},q_{\tau}})$ $=q_{\tau}{\mathcal{J}}^{\textup{track}}({G_{\tau},a^{R+}_{\tau}})$
$+(1-q_{\tau}){\mathcal{J}}^{\textup{track}}({G_{\tau},a^{R-}_{\tau}}).$
@@
In high-complexity object collection tasks, the computed optimal action is to seek assistance if secondary task engagement $G_{t}$ falls below $G_{t}=5$, regardless of other variables. Since low target-tracking performance may be caused by the human paying more attention to the supervisory task, the optimal policy seeks assistance at low target-tracking engagement; thereby allowing the human to finish the tracking task and then assist the manipulator. In contrast, when tracking task engagement is high, the optimal policy depends on human trust: it asks for assistance at low trust levels and collects autonomously for high trust levels. The threshold of trust for seeking assistance decreases with target-tracking performance. Since higher target-tracking performance is an outcome of higher engagement with the tracking task resulting in fewer interruptions by the supervisor, the optimal policy attempts autonomous collection of the object.
@@ IV-C Evaluation Experiment
We recruited $5$ participants, and each one completed two blocks of experiments, during which the robot followed the MPC and greedy policies. The order of the two blocks was randomized. In each block, participants engage in 15 trials each of low-complexity and high-complexity supervisory object collection tasks. Half of these trials involved simultaneous execution of slow-speed target tracking tasks, while the remaining involved normal-speed target tracking. Sequences of object collection and tracking tasks with the above-mentioned complexities were selected and these sequences were randomly permuted for every participant.
@@
To match a realistic scenario in which the participants may not report their trust after each trial, we did not use the reported trust values and estimated trust and tracking task engagement of the human participant using their actions to rely or intervene, and the measured tracking performance $p_{t}$. To this end, we adopted a particle filter with human trust dynamics (1) and engagement dynamics (3), and using the human action model (5) and target-tracking task performance output (4) to estimate $T_{t+1}$ and $G_{t}$.
@@
The cumulative reward for all participants under both the greedy and the MPC policy is illustrated in Fig. 5. The score of a trial is calculated as the sum of rewards in tracking and collection tasks. The median scores are $65.75$ and $57$ for the MPC and the greedy policy, respectively. The MPC policy outperforms the greedy policy for most participants. The cumulative number of interruptions in our MPC policy, $16$, is also lesser than the greedy policy, $23$.
@@
Figure 5: Cumulative reward (sum of collection and tracking rewards) statistics for both policies
V Discussion
When humans encounter a multitasking scenario, they must optimize their resources to achieve acceptable performance in every task. In the context of a human-robot team, human trust in the robot and their engagement in their own responsibilities play a key role in their multi-tasking performance. To understand and leverage human trust and engagement dynamics toward designing better policies for the robot, we developed a model and estimated it using human participant experiment data.
In our prior study [22], we investigated the supervisory object collection task without any secondary task. In contrast to the dual-task paradigm, we observed a higher interruption rate in our prior study (0.173 vs. 0.095). This is possibly due to the division of engagement across two tasks. Additionally, the optimal assistance-seeking policy in the single-task paradigm, as described by [22], seeks assistance only in cases of high task complexity and low human trust. In this paper, the robot’s actions are guided by both the level of human trust and engagement in a secondary task. When engagement with the secondary task is low, resembling a single-task paradigm, the optimal policy consistently seeks assistance. The optimal policy considers the secondary task performance, which tends to decline at low engagement levels, and seeking assistance allowing humans to satisfactorily complete the secondary task before assisting.
Our model parameters indicate that failures decrease trust, aligning with the literature that negative experiences negatively affect trust [11, 13]. Similarly, when humans interrupt and intervene, their trust in the robot tends to decrease [15]. Our human action model suggests that humans are unlikely to rely on robots in scenarios associated with high complexity and low trust. This is consistent with the observations in the literature that at a higher risk, a higher trust is required for humans to rely on the robot [15, 2].
In a dual-task setup, when trust in the robot is low, they tend to interrupt the robot, reducing their attention to the other task and thereby diminishing their performance. When a human supervises an autonomous robot, it has been established that as human trust towards the robot increases, their rate of monitoring and interruption decreases [19]. When performing a dual task, if performance on one task diminishes, it is due to the other task consuming most of the operator’s resources [33]. Therefore, we can consider their engagement in the secondary task as a measure of their trust. When their trust in the robot is low, they tend to supervise it more closely, diverting attention from their tracking task, thus decreasing their engagement and diminishing their performance.
VI Conclusions